OpenStudy (anonymous):

If a charged particle is undeflected while passing through uniform crossed electric and magnetic fields (E and B) where both are mutually perpendicular and perpendicular to the directon of motion, what is the particle's velocity? If the electric field is turned off, what is the charge to mass ratio if it moves in an arc of radius R? I was given no values for this question, so I can assume it only requires the theory. Would someone be able to take me through how I would go about solving this?

OpenStudy (anonymous):

F1 = q E F2 = q v B F1+F2 = 0 qE = - q v B E = - v B v = - E/B circle of radius R : acceleration force = magnetic force (1/2) m v^2 / R = q v B solve for q/m

OpenStudy (anonymous):

Why does F1 + F2 = 0?